OpenStudy (anonymous):

2. M(6, 6) is the midpoint of mc139-1.jpg. The coordinates of S are (8, 9). What are the coordinates of R? (10, 12) (2.5, 4) (7, 7.5) (4, 3) @Ms-Brains

OpenStudy (anonymous):

@Ms-Brains

OpenStudy (poetryprincess):

I am think that it is (4,3)

OpenStudy (poetryprincess):

thinking*

OpenStudy (anonymous):

how did you get that?

MsBrains (ms-brains):

Gimme a sec to solve it.

OpenStudy (anonymous):

okay

MsBrains (ms-brains):

The formula for the midpoint is.. midpoint = ((x + x1)/2, (y + y1)/2) Since M(6,6) is the mipoint, and S(8,9) is the endpoint... (6,6) = ((x + 8)/2, (y + 9)/2) 6 = (x + 8)/2 12 = x + 8 x = 4 6 = (y + 9)/2 12 = y + 9 y = 3 Hence, the coordinates of R is (4,3) in terms of (x,y) Sources: https://answers.yahoo.com/question/index?qid=20110705075101AAluKCd

OpenStudy (anonymous):

oh okay!!! soyou would just find x an y with the formula?